Identified - Our networks team are working on this now, looking into an issue with potential packet loss on the load balancer for these servers that is resulting in intermittent connection issues.
Jul 24, 11:04 BST
Update - Update - This issue looks to be affecting the ability to log in Via FTP, we are continuing to investigate this issue.
Jul 23, 21:35 BST
Update - We are still aware of customers experiencing temporary connection issues on these servers, the sites may have problems displaying content showing a 403 forbidden error or a blank white page.

We are still investigating this and will update once we have further information, we apologise for any ongoing inconvenience caused.
Jul 23, 19:36 BST
Investigating - We are currently seeing abnormal behaviour on these servers in the form of potential connection issues and are investigating the causes.
We expect to have them back to normal operation soon, and apologise for any inconvenience caused.
Jul 23, 14:24 BST
Identified - Following a recent security evaluation of the shared Windows platform we have identified several potential vulnerabilities in the third party application ImageMagick. As the application is deemed vulnerable we will be removing it from all systems with immediate effect. We apologize for any inconvenience this may cause.
Jul 21, 14:53 BST
Investigating - We are currently seeing poor connectivity to wordpress.org, and this is causing problems when attempting to perform automatic updates to WordPress installations.

Our networks team are aware of this and are working on a solution.

We apologise for any inconvenience caused.
Jul 12, 11:22 BST
Investigating - We are seeing that a number of rejections to AOL, this looks to be caused by a limit on their servers restricting the amount of emails that can be received from our shared mailservers. We are currently investigating, our apologies for any inconvenience caused here in the meantime.
Apr 6, 13:04 BST
Update - We are still seeing some rejections and delayed messages to these mail providers and are continuing to work to resolve these problems. We apologise for the ongoing nature of this issue and for the inconvenience caused.
Apr 13, 14:20 BST
Identified - We are seeing that a number of rejections to BT and Yahoo are being caused by badly formed SPF records. If you are having issues sending to addresses with BT and Yahoo from our systems we recommend using the following SPF:-

v=spf1 exists:%{i}.mta.spf.extendcp.co.uk exists:%{i}.web.spf.extendcp.co.uk -all

If you need any assistance in updating this our support team will be happy to help.
Feb 10, 17:52 GMT
Investigating - We are seeing some failures as well as delays with Mail Delivery to BT and Yahoo addresses from messages sent from our network.

Our System Administrators are currently working on this and we will provide further updates here as we have them.

We apologise for the inconvenience caused as a result of this
Jan 23, 12:30 GMT
Monitoring - This is still ongoing, we are continuously monitoring this and will keep this page up to date with the latest information.
Nov 28, 14:21 GMT
Investigating - We are currently seeing a large amount of emails being deferred when sending to hotmail/outlook and other microsoft addresses. This includes and domains that have emails hosted with outlook services.

This looks to be caused by a limit on their servers restricting the amount of emails that can be received from our shared mailservers. Our admin team is currently investigating and attempting to contact them for further details, or a possible work around.

On inspection this may not be effecting all emails sent to these addresses, but a large amount are being blocked. This will cause bouncebacks to the sender, so you can see which emails have been blocked.

Our apologies for any inconvenience caused here in the meantime.
Oct 21, 15:45 BST
Control Panels Operational
DNS Operational
Webservers Degraded Performance
Mailservers Operational
Premium Hosting Operational
VPS & Hybrid Servers Operational
Dedicated Servers Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Past Incidents
Jul 23, 2017
Resolved - This should now be resolved.
Jul 23, 21:09 BST
Monitoring - We believe this is resolved however we are going to continue to monitor this for a couple of hours to be sure.
Jul 23, 19:35 BST
Investigating - We are currently seeing reoccurring abnormal behaviour on this server, and are investigating the causes.

Customers on these servers may experience temporary connectivity issues and with sites offline or not responding, we are aware of this and we hope to have it back to normal operation soon, we apologise for any inconvenience caused.
Jul 23, 18:49 BST
Resolved - This has now been resolved and we will continue to monitor for any unexpected issues.
Jul 23, 18:06 BST
Investigating - We are currently seeing abnormal behaviour on this server, and are investigating the causes.
Customers on these servers may experience temporary connectivity issues and sites offline or not responding, we are aware of this and we hope to have it back to normal operation soon, and apologise for any inconvenience caused.
Jul 23, 17:51 BST
Jul 22, 2017

No incidents reported.

Jul 21, 2017
Resolved - This incident has been resolved.
Jul 21, 08:37 BST
Update - We are still investigating this issue and we apologise for any delay caused regarding this.
Jul 19, 23:20 BST
Investigating - We are currently seeing abnormal behaviour on this servers, and are investigating the causes.
We expect to have it back to normal operation soon, and apologise for any inconvenience caused.
Jul 19, 19:10 BST
Resolved - This incident has been resolved.
Jul 21, 08:37 BST
Investigating - We are still seeing an issue with our domain availability checker and registration tool whereby this fails to query the availability of a specified domain.

This is due to a routing issue between ourselves and our registrar, of which is currently being looked into.

Domain Operations will be affected, we apologise for the inconvenience caused, we expect to have this issue corrected very shortly.
Jul 20, 16:18 BST
Jul 19, 2017
Resolved - This incident has been resolved.
Jul 19, 14:08 BST
Identified - We are currently seeing abnormal behaviour on the following servers,
web57 web58
web59 web60
web67 web68
web69 web70
web189 web190
web191 web192
web298 web299
web300 web301

We are investigating the cause of this and expect to have it back to normal operation shortly.
Jul 19, 09:21 BST
Resolved - This incident has been resolved.
Jul 19, 11:32 BST
Monitoring - Our restore process has finished, and we have re-ran it two more times to be as thorough as possible, but due to the nature of the operation it isn't possible for us to know if every database has been successfully restored. If you find your database is empty please do contact support. Apologies for any inconvenience caused.
Jul 19, 00:23 BST
Investigating - Investigating - We are aware of a problem with the database server that processes the above web-servers. The databases show as empty, or if visiting the site, gives an installation page or connection error.

We are working on restoring the missing databases so that the sites work and this should be completed within the next few hours.

We'll update you as soon as we have more information, we apologise for any inconvenience caused regarding this.
Jul 18, 18:04 BST
Jul 18, 2017
Resolved - This incident has been resolved.
Jul 18, 16:39 BST
Identified - We are still seeing an issue with our domain availability checker and registration tool whereby this fails to query the availability of a specified domain.

This is due to a routing issue between ourselves and our registrar, of which is currently being looked into.

Domain Operations will be affected, we apologise for the inconvenience caused, we expect to have this issue corrected very shortly.
Jul 18, 10:36 BST
Resolved - This incident has been resolved.
Jul 18, 16:39 BST
Update - To resolve this the database server will be switched over. Currently databases are locked into read-only mode to allow us to transfer data over to the new database server. Transfer should be complete within an hour if it continues to progress as it has.
Jul 18, 13:35 BST
Update - Our senior administrators are continuing to look into and work on the problem however we have no timeframe for resolution yet but will update this page as soon as we do.
Jul 18, 10:41 BST
Identified - We are still actually seeing intermittent issues affecting MySQL on the web servers web57, web58, web59 and web60. Our senior administrators are continuing to look into and work on the problem.
Jul 18, 04:13 BST
Monitoring - The issue has been identified and resolved. This is being monitored for any further problems.
Jul 17, 23:41 BST
Investigating - We are currently seeing abnormal behaviour on these webservers, and are investigating the causes.
Customers will experience database connection issues and websites on these servers will be inaccessible.

We expect to have it back to normal operation soon, and apologise for any inconvenience caused.
Jul 17, 18:08 BST
Resolved - This incident has been resolved.
Jul 18, 08:42 BST
Identified - We are currently seeing an issue with our domain availability checker and registration tool whereby this fails to query the availability of a specified domain.

This is due to a routing issue between ourselves and our registrar, of which is currently being looked into.

Domain Operations will be affected, we apologise for the inconvenience caused, we expect to have this issue corrected very shortly.
Jul 17, 09:11 BST
Jul 16, 2017

No incidents reported.

Jul 15, 2017

No incidents reported.

Jul 14, 2017
Resolved - This incident has been resolved.
Jul 14, 12:38 BST
Update - The backup restore has now fully completed. There is a possibility of restore problems for some sites, again as per the last update if you have a website on one of these servers which is still offline please contact our Support Team for further assistance.
Jul 10, 01:34 BST
Monitoring - The majority of affected websites should now be restored. If you have a website on one of these servers which is still offline please contact our Support Team for further assistance.
Jul 9, 12:22 BST
Update - Restores are unfortunately taking longer than expected, however, they are being processed by our system.

We apologise for the continued delay and for the inconvenience caused.
Jul 7, 10:44 BST
Update - The backup restore commenced at 01:30am and we have checked this is running as expected. There is a huge amount of data to restore, which could unfortunately take a few days to fully complete. If you do have your own backups of your files you can of course upload this if you wish, however at some point it may get overwritten by the backup restore currently running. We do sincerely apologise for the inconvenience caused.
Jul 6, 02:02 BST
Update - The new network attached storage device (NAS) has been built and is ready for data to be restored to it. We are currently reviewing the safest and quickest way to do this, and then the restore will be started.
Jul 6, 22:46 BST
Update - Our systems administrators are not able to repair the filesystem, we are working on building the server again and restoring from backups. This will mean minimal data loss, but any changes made since the last backup run will be lost. The majority of websites will be restored from our backups run at approximately 2am this morning, in some cases these may be from an older date.

If you need clarification for this on a per site basis, contact our support team after this has been completed and we can advise when the site was restored from.

We do apologise for any inconvenience caused here.
Jul 5, 22:46 BST
Identified - The cause has been identified as a filesystem error on the network attached storage. Our systems administrators are working on getting this restored.

Further updates will be provided as more information is gathered.
Jul 5, 21:45 BST
Investigating - We are currently having adverse performance on this server, and are investigating the causes.
We expect to have it back to normal operation soon, and apologise for any inconvenience caused.
Jul 5, 21:38 BST
Jul 13, 2017
Completed - The scheduled maintenance has been completed.
Jul 13, 22:29 BST
Scheduled - At 22:00pm (UK time) we will be performing routine maintenance on the Shared Windows environment. During this maintenance servers will be rebooted.
Jul 13, 09:12 BST
Resolved - This incident has been resolved.
Jul 13, 21:43 BST
Investigating - We are currently having adverse performance on this server, and are investigating the causes.
We expect to have it back to normal operation soon, and apologise for any inconvenience caused.
Jul 13, 15:17 BST
Jul 12, 2017
Resolved - This incident has been resolved.
Jul 12, 13:19 BST
Monitoring - We believe the networking issues should be resolved now, but before we mark this status as resolved we will check a few more things. If you do have a problem please contact support.
Jul 12, 07:07 BST
Investigating - We are currently investigating some dedicated servers receiving degraded network performance, where either there is high packet loss or it appears the server is unreachable. This has been raised with our senior network engineers, and further updates will be posted here as soon as we have them. Apologies for any inconvenience caused.
Jul 12, 02:46 BST
Jul 11, 2017
Resolved - 22:00 The high load issue was resolved a few hours previously, and we've been monitoring it since. With no further signs of impact upon the shared platform, this issue can be marked as resolved.
Jul 11, 22:49 BST
Identified - We are currently having adverse performance on our shared hosting platform and are investigating the causes.
We expect to have it back to normal operation soon, and our senior team are working on this as a matter of highest priority and apologise for any inconvenience caused.
Jul 11, 16:46 BST